# Time-varying Graph Signal Estimation via Dynamic Multi-hop Topologies

### v1

@@ V-A Dataset and Preprocessing
Dynamic Multi-hop will be applied to the normal and Autistic subjects selected from the Autism Brain Imaging Data Exchange (ABIDE) dataset, in which the fMRI data was gathered from different internationally shared open databases, 1112 subjects in total [37]. The time-varying node signals $\boldsymbol{x}_{g}[t]$ are provided directly from the data in the form of a multivariate time series, which was formulated by brain regions that are identified from fMRI images. Connectivity within the brain can be measured by transforming neighborhood relationships into a connectivity matrix, where each row and column represents different Regions of Interest (ROI) in the brain [38]. In the experiments, we selected the ROI division of which follows the Harvard-Oxford atlas, which gives us $111$ different regions corresponding to $N_{n}=111$ nodes. The signals within each ROI will naturally form a time-varying node signal $\boldsymbol{x}_{g}[t]$. We selected a normal subject and an Autistic subject within the ABIDE dataset, all with $T=196$ time steps. The noise on the node signals is modeled by Gaussian distribution with zero mean with the variance set corresponding to Signal-to-Noise Ratio (SNR) of 3, 5, and 10.
@@ VII Conclusion and Future Work
Additional improvements can be explored to enhance and extend the Dynamic Multi-hop Model. In the future, we would like to investigate the potential of combining Dynamic Multi-hop with more sophisticated methodologies such as time-series analysis techniques and state-space models. Non-linearity could be introduced into the Dynamic Multi-hop where instead of using graph adaptive filters, we replace the GSP components with GNN components. These extensions will allow Dynamic Multi-hop to make predictions under more complicated scenarios and predict further time steps. Furthermore, the edge pruning based on the SGM can be further improved by more advanced methodologies of edge pruning to produce topologies that can capture the dynamics in the nodes with even higher representation power. Lastly, would also like to extend the fields of applications of our Dynamic Multi-hop to solve additional real-world problems.

### v2

@@ V-A Dataset and Preprocessing
Dynamic Multi-hop will be applied to the normal and Autistic subjects selected from the Autism Brain Imaging Data Exchange (ABIDE) dataset, in which the fMRI data was gathered from different internationally shared open databases [37]. The time-varying node signals $\boldsymbol{x}_{g}[t]$ are provided directly from the data in the form of a multivariate time series, which was formulated by brain regions that are identified from fMRI images. Connectivity within the brain can be measured by transforming neighborhood relationships into a connectivity matrix, where each row and column represents different Regions of Interest (ROI) in the brain [38]. In the experiments, we selected the ROI division which follows the Harvard-Oxford atlas, giving us $111$ different regions corresponding to $N_{n}=111$ nodes. The signals within each ROI will naturally form a time-varying node signal $\boldsymbol{x}_{g}[t]$. We selected a normal subject and an Autistic subject within the ABIDE dataset, all with $T=196$ time steps. The noise on the node signals is modeled by Gaussian distribution with zero mean with the variance set corresponding to Signal-to-Noise Ratio (SNR) of 3, 5, and 10.
@@ VII Conclusion and Future Work
Additional improvements can be explored to enhance and extend the Dynamic Multi-hop Model. Further improvements can be made to enhance and extend the Dynamic Multi-hop Model with advanced methodologies such as time-series analysis techniques and state-space models, which have demonstrated impressive performance both independently and in graph-based formulations [42, 43, 44, 45, 46, 47, 48].” Non-linearity could be introduced into the Dynamic Multi-hop Model by replacing the graph adaptive filters with Graph Neural Network (GNN) components. GNNs are well-suited for capturing complex relationships in graph-structured data and have demonstrated exceptional performance in various machine learning tasks such as classification, regression, and clustering [49]. These extensions will allow Dynamic Multi-hop to make predictions under more complicated scenarios and predict further time steps. Currently, we model the noise using Gaussian distributions. In future work, we aim to evaluate the performance of the Dynamic Multi-hop Model under $\alpha$-stable distributions. As a generalization of Gaussian distributions, $\alpha$-stable distributions are well-suited for capturing impulsive, heavy-tailed behaviors [50, 51]. The robustness of the Dynamic Multi-hop can potentially be improved by the consideration of impulsiveness through $\alpha$-stable noise models. Furthermore, the edge pruning based on the SGM can be further improved by more advanced methodologies of edge pruning to produce topologies that can capture the dynamics in the nodes with even higher representation power. Lastly, would also like to extend the fields of applications of our Dynamic Multi-hop to solve additional real-world problems, such as in material science [42] and image separation [48].